Find the force required to stop a particle of mass 10kg moving with uniform velocity 25m/s .

Grade:12

1 Answers

Karthik
8 years ago
Force=mass*accelaration So,accelaration=0-25/t =-25/t.Force=10*-25/t=-250/t ,you did not provide the time taken for the body to stop